OpenStudy (sbentleyaz):

Let g be a function that is defined for all x, x ≠ 2, such that g(3) = 4 and the derivative of g is g′(x)=${x^2–16}/{x−2}$, with x ≠ 2. Find all values of x where the graph of g has a critical value. For each critical value, state whether the graph of g has a local maximum, local minimum or neither. You must justify your answers with a complete sentence. On what intervals is the graph of g concave down? Justify your answer. Write an equation for the tangent line to the graph of g at the point where x = 3. Does this tangent line lie above or below the graph at this point? Jus

OpenStudy (holsteremission):

\(\bullet\) \(g(x)\) will have critical values wherever \(g'(x)=0\). \(\bullet\) Each critical point can be determined to be the site of a local maximum or minimum by considering the concavity of \(g(x)\) around each critical point. This involves checking the sign of the second derivative \(g''(x)\) in the vicinity of a critical point \(x=c\). \(\bullet\) A description of the function's concavity over the intervals between inflection points can be made by checking the sign of \(g''(x)\) over each interval. \(\bullet\) The line tangent to \(g(x)\) at \(x=3\) will have slope \(g'(3)\). You also know that at \(x=3\), the function takes on the value of \(g(3)=4\). Use the point-slope formula to find the equation of the line itself. \[y-g(3)=g'(3)(x-3)\implies y=\cdots\] \(\bullet\) Checking whether the tangent line lies above or below \(g(x)\) could be done in several ways. The most immediate method might be to simply check the sign of the difference of \(g(x)\) and the tangent line above to either side of their intersection. (e.g. what's the value of \(g(2.9)-y(2.9)\)? of \(g(3.1)-y(3.1)\) ?)

satellite73 (satellite73):

\[ g′(x)=\frac{x^2–16}{x−2}, with x ≠ 2. \]

satellite73 (satellite73):

critical points are where the derivative is zero, i.e. where \(x^2-16=0\) and where it is undefined, i.e. where \(x-2=0\)

satellite73 (satellite73):

to find concavity, you have to find \(g''\) , the derivative of \(\frac{x^2-16}{x-2}\)

OpenStudy (sbentleyaz):

so for critical points would that mean x=2 and x=4?

OpenStudy (sbentleyaz):

@satellite73